# FREE-Merging: Fourier Transform for Model Merging with Lightweight Experts

### v1

@@ Abstract
In the current era of rapid expansion in model scale, there is an increasing availability of open-source model weights for various tasks. However, the capabilities of a single fine-tuned model often fall short of meeting diverse deployment needs. Model merging has thus emerged as a widely focused method for efficiently building a single model tailored for multiple tasks combined from existing models. Nevertheless, existing model merging methods face challenging trade-offs between performance and deployment costs, primarily due to task conflicts within the merged network. Our analysis of neural networks reveals that some task-specific information introduced by fine-tuning minimally enhances performance but heavily impacts generalization, leading to task conflicts. To mitigate the impact of this information, we propose FR-Merging, an innovative method that leverages frequency domain information to efficiently filter harmful specialized information, thereby minimizing the impact of task conflicts on the backbone with minimal cost. Since performance loss is inevitable with cost-free merging methods, we introduce a lightweight task-specific expert that can be dynamically integrated during inference to compensate for information loss. This framework, FREE-Merging (FR-Merging with lightweight experts), strikes a balanced trade-off between training cost, inference speed, storage requirements, and performance. We demonstrate the effectiveness of both FR-Merging and FREE-Merging on multiple tasks across CV, NLP, and Multi-Modal domains and show that they can be flexibly adapted to meet specific needs.
@@ 1 Introduction
With the rapid development of deep learning, the pretrain-finetune paradigm has become the standard methodology [32, 8, 35]. However, in many edge-side applications, users often lack powerful computing resources and large datasets, preferring ready-to-use models tailored to specific scenarios. The recent growth of open-source platforms like HuggingFace [38] has made this more feasible, allowing easy access to model weights for various tasks. Yet, practical demands are often not fully covered by a single existing model and are typically a union of several models [46, 19]. However, if each subtask is handled by different models, it will raise large storage and inference costs in the era of large models. Model merging attempts to directly combine existing model weights to build a model capable of addressing all target tasks [18, 28, 42]. This method mitigates the high training costs and data privacy issues faced by multi-task learning (MTL) [49, 17], reducing both training and deployment costs, and has garnered widespread attention.
@@
However, as model merging seeks to minimize training costs and data, it inevitably encounters performance degradation due to task conflicts [17]. Current solutions fall into three categories. The first method calculates the merging coefficients based on weights or data [19, 28]. The second method uses the parameter differences before and after fine-tuning (task vectors) as the merging target [18, 42, 46]. The third method introduces independent knowledge for each task [17, 24]. Although the third way improves the balance between performance and deployment costs, it faces two issues. On the one hand, it overlooks backbone network optimization during merging, leading to subpar performance. On the other hand, it demands great computation and storage, making it unsuitable for edge deployment.
@@
We believe that developing an effective model merging method requires addressing both of the above issues. On the one hand, since the backbone comprises the majority of parameters and thus determines task performance, an efficient merging mechanism is needed to reduce performance loss. On the other hand, based on the no free lunch theorem, cost-free model merging cannot avoid task conflicts. Introducing task-specific experts is unavoidable. However, it is crucial to ensure that these experts do not significantly burden storage or inference. Therefore, we propose our two-stage optimization method, FREE-Merging, as shown in Fig 1. It provides an efficient backbone and lightweight experts.
@@
Figure 1: The workflow of FREE-Merging involves two main steps. First, FR-Merging utilizes high-pass filtering to remove harmful specialized information from each model, thereby constructing a high-quality merged backbone. Second, it employs lightweight extraction of task experts, which are dynamically added during inference to mitigate the impact of task conflicts.
For the optimization of backbone merging, we propose that fine-tuning specializes parameters for specific tasks [9], so the fine-tuned parameters carry task-specific information. Task-specific information leads to strong task characteristics, which reduces generalization ability and causes task conflicts during merging [45]. Therefore, we aim to mitigate the impact of this information on generalization without affecting model performance, thereby reducing performance loss during merging. Identifying task-specific information in the original parameters has long been challenging [29]. Naturally, the frequency domain comes into our eyes. After the Fourier transform, high-frequency signals represent sharp changes in the signal, while low-frequency components capture the overall structure. Thus, the low-frequency part is more likely to contain the unique aspects of the signal, in this case, the information most relevant to the task. We demonstrate in Sec 4.2 that high-pass filtering of parameters significantly enhances the generalization ability with minimal impact on performance, thus avoiding performance loss due to task conflicts during merging. This computationally lightweight yet highly effective method, FR-Merging, is the first to demonstrate the efficacy of Fourier filtering on neural network parameters, showing potential for broader deep learning applications.
@@
After obtaining a high-performance merged backbone, we utilize a lightweight task-specific expert to compensate for the inevitable information loss during merging. Our goal is to retain task-relevant information with minimal storage. Unlike existing methods that require extensive knowledge storage, our method focuses on rescaling low-frequency signals, which store task-specific information. This allows us to maintain task expertise with only 1% of the parameters. Additionally, to improve inference speed, we further simplify the expert. Inspired by MoE [34], we use a router for dynamic routing of experts, enhancing model flexibility.
@@
In summary, this paper proposes FREE-Merging, which utilizes FR-Merging to enhance model generalization and uses lightweight experts to mitigate information loss of model merging. The main contributions are: 1). We identify the relationship between the low-frequency region of fine-tuning parameters and model generalization, which helps to reduce task conflicts during merging and has significant potential in deep learning. 2). We theoretically validate the necessity of introducing new information in merging and propose an effective lightweight expert construction method to compensate for information loss during merging. 3). According to extensive experiments, the proposed FR-Merging and FREE-Merging demonstrate their effectiveness across various tasks in vision, language, and multi-modal domains.
@@ 2 Related Work
Model Merging. Model merging combines existing model weights to handle multiple tasks without training [18, 19, 43]. The main challenge lies in resolving task conflicts [42]. Simple averaging [39] often leads to significant performance degradation. Methods like Fisher-Merging [28] and RegMean [19] employ different matrices to determine merging coefficients. However, they demand considerable computational costs or access to data, making them unsuitable for edge deployment where both computational power and data are limited. Task Arithmetic [18] merges task vectors instead of weights, while Ties-Merging [42] addresses parameter conflicts, and AdaMerging [46] automates the setup of merging coefficients. Nonetheless, these methods require training and have limited applicability. Techniques such as DARE [47] and PCB-Merging [11] attempt to alleviate performance drops by analyzing parameters, yet they still encounter task conflicts. EMR-Merging [17] and Twin-Merging [24] propose storing knowledge for each task, but they require substantial storage and do not address task conflicts within the merged backbone. Our method uses high-passing filtering to reduce task conflicts in the backbone while utilizing lightweight experts to enhance performance. We achieve a balance between storage, inference speed, and performance, showing impressive results across various tasks. For further details, please refer to Appendix 9.2.
Model Ensemble. Model ensemble combines outputs from multiple models [14, 10], but with large models, it faces storage and inference challenges. In contrast, Model merging enables a single model to perform multiple tasks, greatly reducing storage and inference costs.
Multi-task Learning. Multi-task learning (MTL) aims to solve multiple tasks using a single model [4, 36, 23], but in the era of large models, it faces challenges like high training costs, data privacy issues, and expertise requirements [46]. In contrast, Model merging uses existing open-source model weights to create a multi-task model with little or no training, significantly reducing deployment costs.
@@ 3 Preliminaries
Notation. Assume that $f_{\theta}(x)\rightarrow y$ represents a neural network parameterized by $\theta$, where the input $x\in R^{m}$ and the output $y\in R^{n}$. Recent studies have focused on model merging via task vectors. For task $k$, the task vector is defined as $v_{k}=\theta_{k}-\theta_{pre}$, where $\theta_{k}$ is the fine-tuned model for task $k$, and $\theta_{pre}$ is the corresponding pre-trained model. Additionally, $e_{k}$ represents the task expert that retains the ability to solve task $k$ with few parameters.
Problem Definition. The task of model merging aims to combine models $\{\theta_{k}\}_{k=1}^{K}$ for tasks $T=\{t_{i}\}_{i=1}^{K}$ into a single model $\theta_{m}$ that can handle all tasks in $T$. This paper focuses on a more practical scenario, where a shared pre-trained model $\theta_{pre}$ serves as the base for all fine-tuned models. It is formulated as $\theta_{m}=\lambda\sum_{k=1}^{K}\hat{G}(\theta_{k})$, and when based on the task vector, it becomes $\theta_{m}=\theta_{pre}+\lambda\Sigma_{k=1}^{K}G(v_{k})$, where $\lambda$ is the coefficient, and $\hat{G}/G$ is transformation on the parameters. Assuming a consistent pre-trained model aligns with trends toward foundation architectures, like LLaMa [35] for language and CLIP [32] for vision, which are then fine-tuned for specific tasks.
@@ 4 Methodology
This section presents the proposed FREE-Merging: Sec 4.1 explains the need for a two-part design, Sec 4.2 and Sec 4.3 cover the FR-Merging for the backbone and the lightweight expert module, respectively, followed by the overall FREE-Merging framework and a complexity analysis in Sec 4.5.
@@ 4.1 Motivation
In this section, we discuss why it is necessary to propose an optimized merging method for the backbone and introduce the expert module during inference, respectively.
First, merging optimization for the backbone is crucial, yet often overlooked by current expert-based model merging methods. As the backbone holds most parameters, its performance is key to the overall success. For instance, upgrading from LLaMa2 [35] to LLaMa3 [12] significantly improves results even with the same fine-tuning methods and data. Thus, we propose a training-free backbone merging method for reducing task conflicts in the backbone.
Next, we present a theorem showing the necessity of task experts. Based on Theorem 4.1, introducing additional information is essential. For proof, see Appendix 8, and the effectiveness of task experts is proved in Appendix 10.
@@ 4.2 FR-Merging
In this section, we design an efficient approach to reduce task conflicts in response to the backbone optimization discussed in Sec 4.1. This requires further analysis of the fine-tuning. Fine-tuning can be viewed as mapping the general representation space of a pre-trained model to the space of a specific task [44]. Since different fine-tuning weights fall at different positions in the loss landscape of a task, linearly interpolating the weights could place them in a high-loss region, which is known as task conflict [1]. Therefore, we aim to increase the overlap of the positions of each model weight within the loss landscape of the same task, so that merged models are more likely to fall into a loss basin. Variations in the loss landscape arise from task-specific characteristics in the parameters [40], so we need to eliminate task-specificity in the parameters while preserving the task performance.
If we consider neural network parameters as signals carrying training information, inspired by image processing [31, 3], we transform the task vector into the frequency domain, where high-frequency signals represent sharp changes, and low-frequency components define the overall structure [5]. Since most of the information in the signal is contained within the overall structure, filtering out the low-frequency signals can be seen as removing the most distinctive set of task information. As a result, the likelihood that the linear interpolation falls within the loss basin increases [1]. We use $v(x,y)$ to represent the task vector and $G(x,y)$ to denote the transformed result. Then we have:
@@
$H(\eta,\gamma)$ =$=\begin{cases}1\quad\sqrt{\eta^{2}+\gamma^{2}}\geq D_{0}\\ 0\quad\sqrt{\eta^{2}+\gamma^{2}}<D_{0}\end{cases},$ (2)
where $\mathcal{F}$ and $\mathcal{F}^{-1}$ represent the Fourier transform and inverse Fourier transform, respectively. $D_{0}$ denotes the cutoff frequency, which can be adjusted as a hyper-parameter.
Figure 2: High-pass filtering improves performance over the unfiltered version. Rows represent the added task vectors, and columns indicate the test datasets. The results show a significant boost in generalization with only a slight decline in task performance.
Experimentally, we validate that high-pass filtering significantly improves model generalization with minimal impact on model performance as shown in Fig 2. We fine-tune ViT-B/16 [32] on 30 visual tasks following [17]. After removing low-frequency information, the performance drop on the corresponding tasks (diagonal entries) is substantially smaller than the improvement in generalization (off-diagonal entries). This is a notable finding that similar to image processing, the low-frequency signals in the task vector carry more task-specific information that enhances performance on particular tasks while substantially reducing generalization. Therefore, applying filtering as FR-Merging during model merging can help alleviate task conflicts.
To ensure that the mean of the outputs remains consistent after model merging, we define merging coefficient $\lambda_{i}$ corresponding to the task vector $v_{i}$ for the task $t_{i}$ as:
@@
where $\mathbb{E}$ denotes the average. Overall, we construct a merged backbone in this part that minimizes task conflicts.

### v2

# FREE-Merging: Fourier Transform for Efficient Model Merging

@@ Abstract
With the rapid growth of deep learning, there is an increasing availability of open-source models for various tasks. However, single fine-tuned models often fall short of meeting the diverse needs of users. Model merging has thus emerged as an efficient method to integrate the capabilities of existing models into a unified model. Nevertheless, existing model merging methods face challenging trade-offs between performance and deployment costs, primarily due to task interference. For the first time, we reveal that task interference is evident in the frequency domain of model parameters, yet current efforts only focus on spatial domain solutions, which are largely ineffective in addressing frequency domain interference. To mitigate the impact of frequency domain interference, we propose FR-Merging, an innovative method that effectively filters harmful frequency domain interference on the backbone with minimal computational overhead. Since performance loss is inevitable with cost-free methods, we propose a lightweight task-specific expert module that dynamically compensates for information loss during merging. This proposed framework, FREE-Merging (FR-Merging with experts), strikes a balanced trade-off between training cost, inference latency, storage requirements, and performance. We demonstrate the effectiveness of both FR-Merging and FREE-Merging on multiple tasks across CV, NLP, and Multi-Modal domains and show that they can be flexibly adapted to specific needs.
@@ 1 Introduction
In the current era of deep learning, the pretrain-finetune paradigm has emerged as the standard methodology [37, 9, 40]. However, in numerous edge applications, users often lack powerful computing resources and large datasets, preferring ready-to-use models for specific scenarios. The recent growth of open-source platforms like HuggingFace [43] has made this more feasible, allowing easy access to models for various tasks. Yet, practical demands are often not fully covered by a single existing model and are typically a union of several models [52, 23]. However, running separate models for each subtask increases storage and inference costs, especially with large models. Model merging attempts to combine existing models to build a model capable of addressing all target tasks [21, 33, 48]. This method not only mitigates the training costs and data privacy issues associated with multi-task learning (MTL) [55, 20] but also reduces deployment costs, garnering widespread attention.
@@
While model merging aims to reduce training costs, it often suffers from performance loss due to task interference [20]. To mitigate this issue, existing methods develop along three directions. The first direction computes merging coefficients based on weights or data [23, 33]. The second uses the parameter differences before and after fine-tuning (i.e. task vectors) for merging [21, 48, 52, 57]. The third method introduces independent knowledge for each task [20, 29]. While the third way balances performance and costs, it still faces two critical limitations. First, it neglects backbone optimization during merging, resulting in subpar performance. Second, it requires high computation and storage, making it impractical for edge deployment.
@@
We argue that an effective model merging method requires addressing both of the above issues. First, since the backbone constitutes the majority of parameters and thus determines performance, an efficient backbone merging method is essential to reduce performance loss. Then, given the no free lunch theorem, cost-free merging cannot avoid task interference. Therefore, introducing task-specific experts is necessary. However, they must remain lightweight to avoid storage and inference overhead. Thus, we propose FREE-Merging, a two-stage method that ensures an efficient backbone and lightweight experts as shown in Fig 1.
@@
Figure 1: The workflow of FREE-Merging involves two main steps. First, FR-Merging utilizes high-pass filtering to remove harmful specialized information from each model, thereby constructing a high-quality merged backbone. Second, it employs lightweight extraction of task experts, which are dynamically added during inference to mitigate the impact of task interference.
For backbone merging, the key lies in eliminating task interference [51]. We identify a compelling phenomenon that task interference is evident in the frequency domain of model parameters. And we present in Sec. 4 that frequency-domain interference is strongly correlated with performance. Notably, existing methods, which focus only on spatial-domain operations, fail to address frequency-domain interference, leading to suboptimal outcomes. Experimental evidence reveals that task interference is severe in the low-frequency region. This occurs because, high-frequency signals represent fine-grained variations, while low-frequency parts capture the global structure [18], which is more likely to contain task-specific information that leads to task interference. Leveraging the inherent redundancy in fine-tuned parameters [53, 27, 47], we propose to directly filter out low-frequency parts with severe task interference. As detailed in Sec. 5.2, this high-pass filtering substantially enhances generalization with minimal performance loss, preventing task interference during merging. Our lightweight yet highly effective approach, FR-Merging, is the first to apply Fourier filtering to neural network parameters, showing its potential for broader applications.
@@
After obtaining a high-performance backbone through FR-Merging, we use a lightweight task-specific expert to recover the inevitable information loss with minimal storage. Unlike existing methods that store extensive knowledge, we rescale low-frequency signals, which store task-specific information, retaining task expertise with only 1% of the parameters. Inspired by MoE [39, 56], we employ a router for dynamic expert routing, enhancing model flexibility.
@@
In summary, we introduce FREE-Merging, which utilizes FR-Merging to reduce task interference and incorporates lightweight experts to eliminate information loss of model merging. Our main contributions are: 1). We identify frequency-domain task interference as a critical factor that greatly impacts merging performance but remains challenging to address through spatial-domain methods. 2). To tackle frequency-domain interference, we propose FR-Merging, which reduces interference while preserving performance, constructing the merged backbone. 3). We theoretically validate the necessity of introducing new information during merging and propose an effective expert construction method. 4). Extensive experiments show that FR-Merging and FREE-Merging effectively enhance performance across vision, language, and multi-modal tasks.
@@ 2 Related Work
Model Merging. Model merging integrates existing models to handle multiple tasks without training [21, 23, 49], but it faces task interference challenges [48, 36]. Simple averaging [44] results in great performance degradation. Methods like Fisher-Merging [33] and RegMean [23] determine merging coefficients using various matrices but incur high computational costs or data requirements, limiting their usage in edge deployment. Task Arithmetic [21] merges task vectors instead of weights, while Ties-Merging [48] resolves parameter conflicts, and AdaMerging [52] automates coefficient selection. However, these methods have requirements on data distribution and exhibit limited applicability. Techniques such as DARE [53] and PCB-Merging [12] analyze parameters to mitigate performance drops but still face task conflicts. EMR-Merging [20] and Twin-Merging [29] store task-specific knowledge but require large storage and overlook backbone optimization. Our method applies high-pass filtering to reduce task interference and lightweight experts to enhance performance, achieving a balance between merging costs and effectiveness across various tasks. For further details, please refer to Appendix C.2.
Model Ensemble. Model ensemble combines outputs from multiple models [15, 11, 22], but with large models, it faces storage and inference challenges. In contrast, Model merging enables a single model to solve multiple tasks, significantly reducing storage and inference costs.
Multi-task Learning. Multi-task learning (MTL) aims to solve multiple tasks using a single model [4, 41, 28], but in the era of large models, it faces challenges like high training costs, data privacy issues, and expertise requirements [52]. In contrast, model merging uses existing open-source models to create a multi-task model with little or no training, significantly reducing deployment costs.
@@ 3 Preliminaries
Notation. Assume that $f_{\theta}(x)\rightarrow y$ represents a neural network parameterized by $\theta$, where $x\in R^{m}$ is the input and $y\in R^{n}$ is the output. Recent studies have focused on model merging via task vectors. For task $k$, the task vector is defined as $v_{k}=\theta_{k}-\theta_{pre}$, where $\theta_{k}$ is the fine-tuned model for task $k$, and $\theta_{pre}$ is the corresponding pre-trained model. Additionally, $e_{k}$ represents the task expert that retains the ability to solve the task $k$ with few parameters.
Problem Definition. The objective of model merging is to combine models $\{\theta_{k}\}_{k=1}^{K}$ for tasks $T=\{t_{i}\}_{i=1}^{K}$ into a single model $\theta_{m}$ that can handle all tasks in $T$. We focus on a practical scenario where a shared pre-trained model $\theta_{pre}$ serves as the base for all fine-tuned models. It is formulated as $\theta_{m}=\lambda\sum_{k=1}^{K}\hat{G}(\theta_{k})$, and with task vector, it becomes $\theta_{m}=\theta_{pre}+\lambda\Sigma_{k=1}^{K}G(v_{k})$, where $\lambda$ is the coefficient, and $\hat{G}/G$ is transformation on the parameters. Assuming a shared pre-trained model aligns with trends toward foundation models, like LLaMa [40] for language and CLIP [37] for vision, which are then fine-tuned for specific tasks.
4 Task Interference in Frequency Domain
Figure 2: a) The frequency domain amplitude power distribution of an attention head across different fine-tuned ViT-B-32 models. b) The correlation between the merging performance of ViT-B-32 and the variance of frequency domain amplitude power. Normalized accuracy is defined as the ratio of merged model performance to fine-tuned model performance on the same task.
Task interference is a key factor that leads to the decrease in merging performance. While current methods focus on addressing interference in the spatial domain [48, 50, 53], we observe that task vectors from different fine-tuned models also exhibit great misalignment in the frequency domain.
As shown in Fig. 2(a), we visualize the amplitude power distribution in the frequency domain of task vectors from a ViT-B-32 attention head fine-tuned on various tasks [21]. Here, greater distance from the center represents a higher frequency. Since amplitude indicates signal strength, merging two signals with vastly different strengths inevitably results in the weaker signal being masked. Therefore, task interference in the low-frequency part are severe.
Furthermore, in Fig. 2(b), we train multiple groups of models on eight image tasks [21] and find a clear negative correlation between the variance of amplitude power among models and the normalized merging performance (details can be found in Appendix D.1), confirming the great impact of frequency-domain interference on merging performance.
Method ViT-B/32 ViT-L/14
Task Arithmetic [21] 0.059 0.110
DARE [53] 0.057($\downarrow$3%) 0.108($\downarrow$2%)
Ties-Merging [48] 0.058($\downarrow$2%) 0.109($\downarrow$1%)
Breadcrumbs [7] 0.058($\downarrow$2%) 0.107($\downarrow$3%)
PCB-Merging [12] 0.056($\downarrow$5%) 0.107($\downarrow$3%)
FR-Merging(ours) 0.045($\downarrow$24%) 0.088($\downarrow$20%)
Table 1: The impact of current methods on the mean variance of frequency-domain amplitude across 8 image classification tasks.
Additionally, as shown in Table 1, we show that existing works addressing task interference in the spatial domain fail to effectively mitigate frequency-domain task interference. Therefore, we propose our approach to tackle this issue.
@@ 5 Methodology
This section presents our FREE-Merging: Sec 5.1 explains the need for a two-part design, Sec 5.2 and Sec 5.3 cover the FR-Merging for the backbone and the lightweight expert module, respectively, followed by the FREE-Merging in Sec. 5.4 and a complexity analysis in Sec 5.5.
@@ 5.1 Motivation
We first analyze why both merging optimization for the backbone and the introduction of experts are necessary.
First, backbone optimization is crucial but overlooked by current expert-based merging ways. As the backbone holds most parameters, its performance is key to the overall success. For instance, upgrading from LLaMa2 [40] to LLaMa3 [13] greatly improves results even with the identical fine-tuning. Thus, we propose a training-free merging method to reduce task interference in the backbone.
Next, we present a theorem showing the necessity of task experts. Based on Theorem 5.1, introducing additional information is essential. For proof, see Appendix B, and the effectiveness of task experts is proved in Appendix D.4.
@@ 5.2 FR-Merging
We propose an effective backbone merging method by addressing task interference in the low-frequency fine-tuned parameters as discussed in Sec. 4, which is hard to resolve in the spatial domain. Given the high redundancy of fine-tuned parameters [53], we apply high-pass filtering to directly remove the interfering regions. Let $v(x,y)$ be the task vector and $G(x,y)$ the transformed result. Then we have:
@@
$H(\eta,\gamma)$ =$=\begin{cases}1,\quad\sqrt{\eta^{2}+\gamma^{2}}\geq D_{0}\\ 0,\quad\sqrt{\eta^{2}+\gamma^{2}}<D_{0}\end{cases},$ (2)
where $\mathcal{F}$ and $\mathcal{F}^{-1}$ represent the Fourier and inverse Fourier transform, respectively. $D_{0}$ is the adjustable cutoff frequency. And $\eta$ and $\gamma$ are the distances from the frequency domain center along the $x$- and $y$-axes.
Next, we analyze why high-pass filtering is effective from both intuitive and experimental perspectives. Fine-tuned weights occupy distinct positions in the loss landscape, and linear interpolation often leads to high-loss regions, causing task interference [1, 50]. To address this, we minimize model differences while preserving performance, bringing them closer in the loss landscape to increase the likelihood of the merged model falling into a loss basin [45].
Results in Sec. 4 show great differences in low-frequency regions. This is because, if we consider neural network parameters as signals carrying training information, inspired by image processing [35, 3], high-frequency signals represent sharp variations, while low-frequency parts define the overall structure [5]. Since most information resides in the structure, filtering out low-frequency signals removes key task-specific information. Then, the likelihood that the interpolation falls within the loss basin increases [1].
Figure 3: High-pass filtering improves performance over the unfiltered version. Rows represent the added task vectors, while columns are the test datasets. The results show a significant boost in generalization with only a slight decline in task performance.
Experimentally, we validate that high-pass filtering enhances model generalization with minimal performance loss as shown in Fig 3. We fine-tune ViT-B/16 [37] on 30 visual tasks following [20]. We find that removing low-frequency information results in a smaller performance drop (diagonal) compared to the improvement in generalization (off-diagonal). This indicates that low-frequency parts carry task-specific information that boosts task performance but reduces generalization. Thus, applying filtering like FR-Merging helps alleviate task interference during merging.
To maintain consistent output after merging, we define the merging coefficient $\lambda_{i}$ for the task vector $v_{i}$ as:
@@
where $\mathbb{E}$ represents the average. Overall, we construct a merged backbone in this part that minimizes task conflicts.